OpenStudy (anonymous):

The two sides of a triangle have lengths 10 and 18. Which inequalities describe the values that possible lengths for the third side?

OpenStudy (anonymous):

\[a) 8 \le x \le 28 or b) 8 < x < 28 c) 10 < x < 18 d) 10 \le x \le 18\]

OpenStudy (anonymous):

those are our options

OpenStudy (anonymous):

Let x equal the third side. \[10 + 18 \ge x\]

OpenStudy (anonymous):

0<x<10+18 (x can't be 0)

OpenStudy (anonymous):

and it also can't be equal to 10+18, that would be a straight line

OpenStudy (anonymous):

so It's not C?

OpenStudy (anonymous):

This is my first time doing a problem like this, so I need a clear explanation

OpenStudy (anonymous):

you can call it "c", if you want... "x" is used for any unknown number